OpenStudy (anonymous):

5. A wooded area in a state park has a mixture of different types of trees. There are 800 pine trees and 50 oak trees. The number of pine trees is decreasing at a rate of 5% per year. The number of oak trees is increasing at a rate of 15% per year. If these trends continue: A. Write two functions to model this situation and graph those two functions on the same coordinate grid.

OpenStudy (anonymous):

I think its f(x)=800-800(0.05) and then f(x)= 50- 50(0.15)

OpenStudy (anonymous):

but I don't think that's right... can someone please help

OpenStudy (anonymous):

The number of oak trees is increasing, so the function should produce more trees over time.

OpenStudy (anonymous):

so is it f(x)=50+50(0.15)

OpenStudy (anonymous):

@jN7

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

is my first equation for the other tree right?

OpenStudy (anonymous):

basically it would be f(x)=50+(50*.15x)

OpenStudy (anonymous):

oh 0k

OpenStudy (irishboy123):

you need to include compounding. at the minute you are adding or deducting x% of the "original" amount from the current balance. you need to add or deduct x% of the current amount from current. does that make sense?

OpenStudy (anonymous):

yes it does... thank you